OpenStudy (anonymous):

Write the equation of the line that passes through the points (-1, -4) and (2, 5). Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did. (2 points)

OpenStudy (anonymous):

slope=m=(y2-y1)/(x2-x1)=(5-(-4))/(2-(-1))=9/3=3 y=mx+b=3x+b choose (-1,-4) --> -4=3(-1)+b --> b=-1 Therefore, y=3x-1

OpenStudy (anonymous):

It does not matter which point you choose to substitute in the linear equation to find the value b, since the a line with a given slope (m) that is specified by two points is unique and no other lines with a different slope can pass through the two points.

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

No problemo..